OpenStudy (anonymous):

What is the value of y in the solution to the system of equations? 1/3x + 1/4y = 1 2x – 3y = –30 a.–8 b.–3 c.3 d.8

OpenStudy (anonymous):

@airjordan200

OpenStudy (anonymous):

help?

OpenStudy (mathmale):

In typing the fractions 1/3 and 1/4, you have inadvertently introduced ambiguity. How would your reader know whether you meant (1/3) x or 1 / (3x) ? Please use parentheses to eliminate that ambiguity.

OpenStudy (mathstudent55):

Since the first equation has denominators 3 and 4, multiply the entire equation by the LCM of the two denominators. That will eliminate the denominators. Can you do that and show your work?

OpenStudy (mathstudent55):

What is the least common multiple of 3 and 4?

OpenStudy (anonymous):

12?

OpenStudy (mathmale):

1/3x + 1/4y = 1 2x – 3y = –30 By asking "what is the least common multiple of 3 and 4," mathstudent55 is assuming that you mean (1/3)x + (1/4)y. It takes just a second or two longer: eliminate ambiguity by using parentheses to indicate what operations are to be done first. Once you've eliminated the fractions from the first equation, we can move on to solve the system of linear equations for y as required by the problem statement.

OpenStudy (anonymous):

i know but i get like 3.666666666 continuing

OpenStudy (anonymous):

and i have like 2 min left so...

OpenStudy (mathstudent55):

Yes. 12 is the LCM. Now multiply the entire first equation by 12.

OpenStudy (mathmale):

Sorry, but I don't understand how 3.66666.... is related to asking you to use parentheses. Like mathstudent55, I'm going to assume that you mean (1/3) x + (1/4)y = 1. Yes, the LCD is 12. Clear this equation of fractions, pleasae.

OpenStudy (mathmale):

4x + ?y =12

OpenStudy (anonymous):

3?

OpenStudy (mathmale):

Yes. 12 times (1/4) is 3. Yes. 4x + 3y = 12 and the second equation is 2x – 3y = –30 How would you now solve for y?

OpenStudy (mathstudent55):

This is how you multiply the entire first equation by 12. You multiply each side by 12. On the left side, you then distribute the 12. Can you simplify it now? What is 12 * 1/3 ? What is 12 * 1/4 ?